OpenStudy (anonymous):

solve: -5y+3x=-16 10y+4x=62

OpenStudy (cwrw238):

multiply ist equn. by 2 -10y+6x = -32 10y + 4x = 62 add: 10x = 30 x=3 plug this into ist equn.: -5y + 9 = -16 -5y = -25 y = 5 x=3,y=5

OpenStudy (anonymous):

how did u get 30?